For Kepler's period law, what is the relation if T and R is respectively the period of revolution of the planet about the sun and distance of the planet from the sun?
1
Expert's answer
2020-02-17T09:19:07-0500

For Kepler's period law


"\frac{T^2}{R^3}=\frac{4\pi^2}{GM_S}"

"\frac{T^2}{R^3}=\frac{4\pi^2}{(6.67\cdot 10^{-11})(1.99\cdot 10^{30})}=2.97\cdot 10^{-19}\frac{s^2}{m^3}"